OpenStudy (shelby1290):

A football is punted into the air. Its height, h, in metres, after t seconds is given by h(t)=1+24.5t-4.9t^2 a) Find the maximum height of the ball b) When does the ball reach its maximum height?

OpenStudy (anonymous):

is this flvs

OpenStudy (welshfella):

you can find the maximum height by converting the equation to vertex form or by using calculus

OpenStudy (shelby1290):

@DaBest21 no

OpenStudy (shelby1290):

@welshfella do i have to complete the square?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

h(t)=1+24.5t-4.9t^2 = -4.9(t^2 - 5t) + 1 = -4.9[(t - 2.5)^2 - 2.5^2] + 1

OpenStudy (welshfella):

maximum height is when t = 2.5 h(2.5) = -4.9 * -6.25 + 1

OpenStudy (welshfella):

time to reach maxm height is 2.5 seconds maxm height = -4.9*-6.25 + 1

OpenStudy (welshfella):

ok?

OpenStudy (shelby1290):

@welshfella how did you get 2.5 seconds?

OpenStudy (mathmale):

Shelby, have you been studying calculus, or is this problem from an algebra course?

OpenStudy (welshfella):

5 /2 = 2.5 - completing the square

OpenStudy (shelby1290):

@mathmale no i'm not studying calculus. I'm taking Functions but this question is based on quadratics

OpenStudy (welshfella):

i'll leave you in capable hands of Mathmale

OpenStudy (mathmale):

OK, then your welshfella is correct in seeking the vertex of the curve by completing the square. However, there are other ways to do that. You could re-write the equation in the form ax^2+bx+c and then find the x-coordinate of the vertex by calculating -b/2a.

OpenStudy (mathmale):

That is, t=-b/2a. This t value represents the number of seconds it took the ball to reach its max height.

OpenStudy (shelby1290):

alright, i'll try writing this out. one sec

OpenStudy (mathmale):

Given h(t)=1+24.5t-4.9t^2, identify the coefficient of the t^2 term. This is "a." Then identify the coeff. of the t term. This is "b"

OpenStudy (shelby1290):